OpenStudy (anonymous):

what is the value of x/y if 28/x^2=7/xy? F. 4 G. 4x^2 H. 4xy J. 28y K. 196

OpenStudy (anonymous):

Solve the following for x:\[\frac{28}{x^2}\text{=}\frac{7}{x y}\]\[x=4 y \]Can you now guess what x/y is?

hero (hero):

\[\frac{28}{7} = \frac{x^2}{xy}\] \[\frac{7}{7}\left(\frac{4}{1}\right) = \frac{x}{x}\left(\frac{x}{y}\right)\] \[\frac{4}{1} = \frac{x}{y}\]